# A linear Convergence result for the Jacobi-Proximal Alternating Direction Method of Multipliers

### v2

@@ 1.2. Contribution
The main contribution in this paper are as follows. We establish the global linear convergence of Jacobi-Proximal ADMM(Algorithm 1) under the assumption that the objective function $f_{i}$ are strictly convex and $L_{i}$-smooth. We establish its convergence at a rate of $O(\sigma^{k})$ for the optimality measure defined in (2.4). Our numerical results on the LCQP model and the optimal resource allocation problem support the linear convergence result of this work.
1.3. Organization
We organize the rest of this paper as follows. In Section 2, we present a sufficient condition to ensure the convergence of (2.2). Then based on the analysis in Section 2, we prove the global convergence of the Algorithm 1 in Section 3. In Section 4, we present convergence results by conducting numerical experiments on two examples. Finally, we conclude the paper in Section 5.
2. Statement of the main result
@@
In this section, we state the main result of this paper concerning the linear convergence of the Jacobi-Proximal ADMM in Algorithm 1. Since both of the objective function and constraints of (1.3) are sums of terms on individual $x_{i}$’s, the problem can be decomposed into $N$ smaller subproblems, which is solved in a parallel and distributed manner.
@@
We state assumptions on the cost functions which will be used throughout the paper. Here we let $n=n_{1}+\cdots+n_{N}$ and set $A\in\mathbb{R}^{m\times n}$ by $A=\big{[}A_{1}\cdots A_{N}\big{]}$. Also, we set $x^{k}\in\mathbb{R}^{n}$ for $k\geq 0$ and $x^{*}\in\mathbb{R}^{n}$ as $x^{k}=\big{[}x_{1}^{k};~{}\cdots;~{}x_{N}^{k}\big{]}$ and $x^{*}=\big{[}x_{1}^{*};~{}\cdots;~{}x_{N}^{*}\big{]}$. We then have
@@
We assume that the matrix $\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}\in\mathbb{R}^{n\times m}$ with $n=n_{1}+\cdots+n_{N}$ has full column rank.
Here $P_{i}$ is positive semi-definite for each $1\leq i\leq N$. For $k\geq 0$, we denote by $\left(x_{1}^{k},\cdots,x_{N}^{k};\lambda^{k}\right)$ the variable generated by the Jacobi-Proximal ADMM. We state the main result of this paper.
@@
Suppose Assumptions 2.1-2.5 hold. Choose any $s>0$ such that
@@
and suppose that $20<\gamma<$ 2, $\rho>0$ are chosen so that there exist values $\xi_{i}>0$ such that
@@
Let $\sigma=\max\left\{{1-2\gamma\rho sc_{A}^{2},\mu_{s}}\right\}$. Here $c_{A}>0$ and $\mu_{s}\in(0,1)$ are constants satisfying
[…]
where the first inequality ensures that $c_{A}=\sqrt{\min\textrm{eig}\Big{(}\sum_{i=1}^{N}A_{i}A_{i}^{T}\Big{)}}>0$. Then we have
@@
where $\phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is defined as $\phi\big{(}x^{k},\lambda^{k}\big{)}=\frac{1}{2\gamma\rho}\big{\|}\lambda^{k}-\lambda^{*}\big{\|}^{2}+\frac{1}{2}\sum_{i=1}^{N}\big{\|}x_{i}^{k}-x_{i}^{*}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}$. As a result of (2.4), we have the following result
@@
3. Convergence
In this section, we prove Theorem 2.6. For this, we prepare the several lemmas.
@@ 3. Convergence
$c_{A}=\min_{\lambda\in\mathbb{R}^{m}\backslash\{0\}}\frac{\Big{(}\sum_{i=1}^{N}\|A_{i}^{T}\lambda\|^{2}\Big{)}^{\frac{1}{2}}}{\|\lambda\|}=\min_{\lambda\in S^{m-1}}\frac{\Big{(}\sum_{i=1}^{N}\|A_{i}^{T}\lambda\|^{2}\Big{)}^{\frac{1}{2}}}{\|\lambda\|}.$
@@
Thus, we obtain the final result for $\mu_{s}\in(0,1)$, which is sufficiently close to $1$. ∎
@@
If one chooses some $\xi_{i}>0$ such that $20<\gamma<$ 2, $\rho>0$, and $s>0$ satisfies the following conditions:
@@
for $\xi_{j}>0~{}(1\leq j\leq N)$. Adding $\sum_{i=1}^{N}\big{\|}x_{i}^{k+1}-x_{i}^{k}\big{\|}_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}$, $-8s\sum_{i=1}^{N}\bigg{\|}\Big{(}\rho A_{i}^{T}A_{i}+P_{i}\Big{)}\big{(}x_{i}^{k+1}-x_{i}^{k}\big{)}\bigg{\|}^{2}$, and $\frac{2-\gamma}{\gamma^{2}\rho}\big{\|}\lambda^{k+1}-\lambda^{k}\big{\|}^{2}$ to both sides of (3.15), we obtain
@@
4. Numerical experiments
In this section, we present numerical tests of Algorithm 1 supporting the linear convergence property proved in Theorem 2.6. We test the Algorithm 1 for the linearly const rained quadratic programming (LCQP) model [34] and the optimal source allocation problem [37].
4.1. The LCQP Model
@@
where $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}~{}\textrm{is given by}~{}f_{i}(x_{i})=\frac{1}{2}x_{i}^{T}H_{i}x_{i}+x_{i}^{T}q_{i}\textrm{ with}~{}H_{i}\in\mathbb{S}_{+}^{n_{i}},~{}q_{i}\in\mathbb{R}^{n_{i}},~{}A_{i}\in\mathbb{R}^{m\times n_{i}}~{}\textrm{and}~{}c\in\mathbb{R}^{m}$. Since $H_{i}\in\mathbb{S}_{+}^{n_{i}}$, the quadratic function $f_{i}$ is strongly convex for each $i=1,\ldots,N$. We take $n_{1}=\cdots=n_{N}:=n$ in our experiment. We conduct experiments with the damping parameter $\gamma\in\{0.5,1.0,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.1,1,5,10\}$ in Algorithm 1. The matrices involved in this problem are randomly generated through the following procedure:
  1. (1)
@@
Figure 1. Experimental results on 3-block LCQP model for fixed $\gamma$ under varying $\rho$.
Figure 2. Experimental results on 3-block LCQP model for fixed $\rho$ under varying $\gamma$.
Figure 3. Experimental results on 10-block LCQP model for fixed $\gamma$ under varying $\rho$.
Figure 4. Experimental results on 10-block LCQP model for fixed $\rho$ under varying $\gamma$.
4.2. The Optimal Resource Allocation Problem
@@
where $f_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is given as $f_{i}(x_{i})=(1/2)a_{i}(x_{i}-c_{i})^{2}+\log[1+exp(b_{i}(x_{i}-d_{i}))]$ with the coefficients $a_{i},b_{i},c_{i},d_{i}$ generated randomly with uniform distributions on $[0,2]$, $[-2,2]$, $[-10,10]$, $[-10,10]$ respectively. We conduct experiments with the damping parameter $\gamma\in\{0.5,1.0,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.1,1,5,10\}$ in Algorithm 1. We update $x_{i}^{k+1}$ in (1.5) using the equation derived from the optimality given as follows:
@@
In order to study the convergence behavior of the test algorithms, we test two cases: $6$-block and $20$-block. The value $u^{*}=\big{(}x_{1}^{*},\ldots,x_{N}^{*};\lambda^{*}\big{)}$ is obtained numerically by running Algorithm 1 $5,000$ iterations. We employ the same error metric $\text{dis}(u^{k})$ as defined in (4.2) for this experiment. The initial values $u^{0}=(x_{1}^{0},\ldots,x_{N}^{0};\lambda^{0})$ are set to zero. The experimental results are plotted in Figures 5 and 6 for the $6$-block case, and in Figures 7 and 8 for the $20$-block case. We observe that the value $\text{dis}(u^{k})$ converges linearly to zero, as anticipated in Theorem 2.6.
Figure 5. Experimental results on 6-block problem for the optimal resource allocation model with a fixed $\gamma$ under varying $\rho$.
Figure 6. Experimental results on 6-block problem for the optimal resource allocation model with a fixed $\rho$ under varying $\gamma$.
Figure 7. Experimental results on 20-block problem for the optimal resource allocation model with a fixed $\gamma$ under varying $\rho$.
Figure 8. Experimental results on 20-block problem for the optimal resource allocation model with a fixed $\rho$ under varying $\gamma$.
5. Conclusion
In this paper, we consider a multi-block Jacobi-Proximal ADMM algorithm for solving linearly coupled optimization problems in a parallel way. We established the linear convergence of the algorithm and derived its convergence rate involving more than two variable blocks when the cost functions are strongly convex and Lipschitz continuous. This is the first linear convergence result for the Jacobi-Proximal ADMM.

### v4

@@ 1.2 Contribution
The main contribution in this paper are as follows. We establish a global linear convergence of Jacobi-Proximal ADMM(Algorithm 1) under the assumption that the objective functions $f_{i}$ are strictly convex and $L_{i}$-smooth. Our numerical results on the LCQP model and the optimal resource allocation problem support the linear convergence result of this work.
1.3 Organization
We organize the rest of this paper as follows. In Section 2, we state the assumptions and the main result of this paper. Section 3 is devoted to prove the main result. In Section 4, we present numerical experiments supporting the linear convergence result.
2 Statement of the main result
@@
In this section, we state the main result of this paper concerning the linear convergence of the Jacobi-Proximal ADMM described in Algorithm 1. Since both of the objective function and constraints of (1.1) are sums of terms on individual $x_{i}$’s, the problem can be decomposed into $N$ smaller subproblems, which is solved in a parallel and distributed manner.
@@
Here we let $n=\sum_{i=1}^{N}n_{i}$ and set $A\in\mathbb{R}^{m\times n}$ by $A=\big[A_{1}\cdots A_{N}\big]$. For $k\geq 0$, we denote by $\left(x_{1}^{k},\cdots,x_{N}^{k};\lambda^{k}\right)$ the variable generated by Algorithm 1. Also, we set $x^{k}\in\mathbb{R}^{n}$ for $k\geq 0$ and $x^{*}\in\mathbb{R}^{n}$ as $x^{k}=\big[x_{1}^{k};\cdots;x_{N}^{k}\big]$ and $x^{*}=\big[x_{1}^{*};\cdots;x_{N}^{*}\big]$. With this notation, we have
@@
We use $\|z\|$ to denote the standard Euclidean norm of $z$, and $\langle\cdot,\cdot\rangle$ to denote the standard inner product. For a positive definite matrix $M\in\mathbb{R}^{l\times l}$, we define the $M$-norm as $\|z\|_{M}:=\sqrt{z^{T}Mz},\quad\forall z\in\mathbb{R}^{l}$. If the matrix $M$ is positive semi-definite, then $\|\cdot\|_{M}$ is a semi-norm. We state assumptions on the cost functions which will be used throughout the paper.
@@
We assume that there exists a minimizer of problem (1.1) satisfying the KKT condition as follows.
@@
We assume that the matrix $\big[A_{1}^{T};~A_{2}^{T};~\cdots;~A_{N}^{T}\big]\in\mathbb{R}^{n\times m}$ with $n=\sum_{i=1}^{N}n_{i}$ has full column rank.
We state the main result of this paper.
@@
Suppose Assumptions 2.1-2.5 hold and assume that $P_{i}$ is a positive semi-definite matrix for each $1\leq i\leq N$. Choose any $s>0$ such that
@@
and suppose that $0<\gamma<2$, $\rho>0$ are chosen so that there exist values $\xi_{i}>0$ such that
@@
Define $\sigma=\max\left\{{1-2\gamma\rho sc_{A}^{2},\mu_{s}}\right\}\in(0,1)$ for any $c_{A}\in\Big(0,\frac{1}{\sqrt{2\gamma\rho s}}\Big)$, and $\mu_{s}\in(0,1)$ satisfying
$\left\{\begin{aligned} &c_{A}^{2}I_{m}\preceq\sum_{i=1}^{N}A_{i}A_{i}^{T},\\ &(\rho+4Ns\rho^{2}D)A_{i}^{T}A_{i}+P_{i}\preceq\mu_{s}(\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-Ls)I_{n_{i}}.\end{aligned}\right$.
Then we have
@@
where $\phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is defined as
$\phi\big(x^{k},\lambda^{k}\big)=\frac{1}{2\gamma\rho}\big\|\lambda^{k}-\lambda^{*}\big\|^{2}+\frac{1}{2}\sum_{i=1}^{N}\big\|x_{i}^{k}-x_{i}^{*}\big\|_{\rho A_{i}^{T}A_{i}+P_{i}+2(\alpha-2Ls)I_{n_{i}}}^{2}.$
As a result of (2.4), we have the following linear convergence result
@@
3 Proof of the main result
In this section, we give the proof of Theorem 2.6. For this, we prepare several lemmas.
@@ 3 Proof of the main result
[…]
@@
Thus, we obtain the final result for $\mu_{s}\in(0,1)$, which is sufficiently close to 11. ∎
@@
If one chooses some $\xi_{i}>0$ such that $0<\gamma<2$, $\rho>0$, and $s>0$ satisfies the following conditions:
@@
for $\xi_{j}>0~(1\leq j\leq N)$. Adding $\sum_{i=1}^{N}\big\|x_{i}^{k+1}-x_{i}^{k}\big\|_{\rho A_{i}^{T}A_{i}+P_{i}}^{2}$, $8s\sum_{i=1}^{N}\bigg\|\Big(\rho A_{i}^{T}A_{i}+P_{i}\Big)\big(x_{i}^{k+1}-x_{i}^{k}\big)\bigg\|^{2}$, and $\frac{2-\gamma}{\gamma^{2}\rho}\big\|\lambda^{k+1}-\lambda^{k}\big\|^{2}$ to both sides of (3.15), we obtain
@@
4 Numerical experiments
In this section, we present numerical tests of Algorithm 1 supporting the linear convergence property proved in Theorem 2.6. We test the Algorithm 1 for the linearly const rained quadratic programming (LCQP) model [A_faster] and the optimal source allocation problem [Optimal_scaling]. In Section 4, all numerical experiments were performed using Python 3.10.16 with NumPy 2.1.3, SciPy 1.10.1, and Matplotlib 3.9.2 on a workstation with an Intel Core i9-12900KS CPU (3.40 GHz) and 32.0 GB RAM.
4.1 The LCQP Model
@@
where $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}~\textrm{is given by}~f_{i}(x_{i})=\frac{1}{2}x_{i}^{T}H_{i}x_{i}+x_{i}^{T}q_{i}\textrm{ with}~H_{i}\in\mathbb{S}_{+}^{n_{i}},~q_{i}\in\mathbb{R}^{n_{i}},~A_{i}\in\mathbb{R}^{m\times n_{i}}~\textrm{and}~c\in\mathbb{R}^{m}$. Since $H_{i}\in\mathbb{S}_{+}^{n_{i}}$, the quadratic function $f_{i}$ is strongly convex for each $i=1,\ldots,N$. We take $n_{1}=\cdots=n_{N}:=n$ in our experiment. We conduct experiments with the damping parameter $\gamma\in\{0.1,0.5,1.5,1.9\}$, and the penalty parameter $\rho$, chosen as $\{0.03,1,5,10\}$ for $N=3$ and $\{10^{-5},0.1,5,10\}$ for $N=10$, in Algorithm 1. The matrices involved in this problem are randomly generated through the following procedure:
  1. 1.
@@
As shown in Figure 1, for fixed values of the damping parameter $\gamma\in\{0.1,0.5\}$, the algorithm exhibits the fastest convergence when $\rho=1$. This indicates that a smaller penalty parameter does not necessarily lead to faster convergence, emphasizing the importance of proper choice of the penalty parameter $\rho$ for the fast convergence. Figure 2 shows the effect of the damping parameter $\gamma$ under fixed $\rho$. The convergence rate improves as $\gamma$ increases when $\rho=0.03$, whereas for $\rho=5$, the algorithm exhibits similar convergence rates across all tested values of $\gamma$, indicating limited sensitivity to the damping parameter.
Figure 1: Experimental results on LCQP model for $N=3$ ($m=100$, $n=40$), with fixed $\gamma$ and varying $\rho$.
Figure 2: Experimental results on LCQP model for $N=3$ ($m=100$, $n=40$), with fixed $\rho$ and varying $\gamma$.
Figure 3: Experimental results on LCQP model for $N=10$ ($m=100$, $n=60$), with fixed $\gamma$ and varying $\rho$.
Figure 4: Experimental results on LCQP model for $N=10$ ($m=100$, $n=60$), with fixed $\rho$ and varying $\gamma$.
4.2 The Optimal Resource Allocation Problem
@@
where $f_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is given as $f_{i}(x_{i})=(1/2)a_{i}(x_{i}-c_{i})^{2}+\log[1+exp(b_{i}(x_{i}-d_{i}))]$ with the coefficients $a_{i},b_{i},c_{i},d_{i}$ generated randomly with uniform distributions on $[0,2]$, [2,2][-2,2], [10,10][-10,10], [10,10][-10,10] respectively. We conduct experiments with the damping parameter $\gamma\in\{0.1,0.5,1.5,1.9\}$, and the penalty parameter $\rho\in\{0.03,1,5,10\}$ in Algorithm 1. We update $x_{i}^{k+1}$ in (1.5) using the equation derived from the optimality given as follows:
@@
In order to study the convergence behavior of the test algorithms, we test two cases: $6$-block and $20$-block. The value $u^{*}=\big(x_{1}^{*},\ldots,x_{N}^{*};\lambda^{*}\big)$ is obtained numerically by running Algorithm 1 $4,000$ iterations. We employ the same error metric $\text{dis}(u^{k})$ as defined in (4.2) for this experiment. The initial values $u^{0}=(x_{1}^{0},\ldots,x_{N}^{0};\lambda^{0})$ are set to zero. The experimental results are plotted in Figures 5 and 6 for the $6$-block case, and in Figures 7 and 8 for the $20$-block case. We observe that the value $\text{dis}(u^{k})$ converges linearly to zero, as anticipated in Theorem 2.6.
Figure 5 illustrates that, for fixed damping parameters $\gamma\in\{0.1,0.5,1.5\}$, the algorithm achieves the fastest convergence when $\rho=1$. This suggests that simply choosing a smaller penalty parameter does not guarantee faster convergence, highlighting the need for proper choice of $\rho$ for the fast convergence. In Figure 6, the effect of the damping parameter $\gamma$ is investigated at fixed value of $\rho$. When $\rho=0.03$, increasing $\gamma$ leads to improved convergence rates. In contrast, for $\rho=5$, the convergence rates remain similar regardless of the value of $\gamma$, implying that the algorithm’s performance is relatively insensitive to changes in $\gamma$.
Figure 5: Experimental results on the optimal resource allocation model for $N=6$, with fixed $\gamma$ and varying $\rho$.
Figure 6: Experimental results on the optimal resource allocation model for $N=6$, with fixed $\rho$ under varying $\gamma$.
Figure 7: Experimental results on the optimal resource allocation model for $N=20$, with fixed $\gamma$ under varying $\rho$.
Figure 8: Experimental results on the optimal resource allocation model for $N=20$, with fixed $\rho$ under varying $\gamma$.
5 Conclusion
In this paper, we study the Jacobi-Proximal ADMM algorithm designed to solve the linearly coupled optimization problems in a parallel way. We achieved the first result for the linear convergence of the algorithm and derived its convergence rate when the cost functions are strongly convex and Lipschitz continuous. This is the first linear convergence result for the Jacobi-Proximal ADMM. We also presented numerical experiments supporting the convergence result.
Declarations
Funding
No funding was received for conducting this study.
Competing interests
The authors have no relevant financial or non-financial interests to disclose.
Code availability
The Python code used for the numerical experiments in Section 4 is publicly available at https://github.com/hyelinchoi/Jacobi_Proximal_ADMM.